OpenStudy (anonymous):

Which of the following are solutions to the equation below? (4x - 1)2 = 11

OpenStudy (solomonzelman):

\(\Large\color{green}{ \bf (4x - 1)^2 = 11}\) Do you prefer to complete the square, do a quadratic formula, or factor ?

OpenStudy (anonymous):

Quadratic Formula sir

OpenStudy (anonymous):

or ma'am

OpenStudy (solomonzelman):

just a teenager, if you want, a sir :) ..... \(\Large\color{green}{ \bf (4x - 1)^2 = 11}\) I am expanding the left side... (using the rule: (a-b)^2=a^2-2ab+b^2 ) \(\Large\color{green}{ \bf 16x^2-8x+1 = 11}\) now I am going to subtract 11 from both sides. \(\Large\color{green}{ \bf 16x^2-8x-10 = 0}\)

OpenStudy (solomonzelman):

just a teenager, if you want, a sir :) ..... \(\Large\color{green}{ \bf (4x - 1)^2 = 11}\) I am expanding the left side... (using the rule: (a-b)^2=a^2-2ab+b^2 ) \(\Large\color{green}{ \bf 16x^2-8x+1 = 11}\) now I am going to subtract 11 from both sides. \(\Large\color{green}{ \bf 16x^2-8x-10 = 0}\)

OpenStudy (solomonzelman):

I got disconnected -:(

OpenStudy (anonymous):

It's okay

OpenStudy (solomonzelman):

\(\Large\color{brown}{ \bf 16x^2-8x-10 = 0}\) I am going to factor out of 2... \(\Large\color{brown}{ \bf 2(8x^2-4x-5) = 0}\) now I am going to divide both sides by 2, which will just allow me to get rid of the 2 on the left side. \(\Large\color{brown}{ \bf 8x^2-4x-5 = 0}\) Good so far ?

OpenStudy (anonymous):

Think so

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \bf \frac{-b±\sqrt{b^2-4ac} }{2a} }\)

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \bf \frac{-(-4)±\sqrt{(-4)^2-4(8)(-5)} }{2(8)} }\)

OpenStudy (anonymous):